OpenStudy (anonymous):

Find the compound inequality 7>-4x+6 or 10≤-4x+3

OpenStudy (anonymous):

4>-4x+3 or 10≤-4x+3

OpenStudy (anonymous):

-4x+3 < 4or 10≤-4x+3

OpenStudy (anonymous):

10≤-4x+3< 4

OpenStudy (anonymous):

tht is the answer

OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

7>-4x+6 or 10≤-4x+3 not 4>-4x+3 or 10≤-4x+3\[\text{Reduce}[7>-4x+6 \| 10\leq -4x+3] =\left( x\leq -\frac{7}{4}\|x>-\frac{1}{4}\right) \]